Coney Island Entertainment issues $1,500,000 of 6% bonds, due in 10 years, with interest payable semiannually on June 30 and December 31 each year.

   

Calculate the issue price of a bond and complete the first three rows of an amortization schedule when:

Required:

1. The market interest rate is 6% and the bonds issue at face amount. (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use appropriate factor(s) from the tables provided. Do not round interest rate factors.)

2. The market interest rate is 7% and the bonds issued at a discount. (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use appropriate factor(s) from the tables provided. Do not round interest rate factors.)

3. The market interest rate is 5% and the bonds issued at a premium. (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use appropriate factor(s) from the tables provided. Do not round interest rate factors.)
Answer #1

Solution 1:

Chart Values are based on:
n= (10 Years*2) 20 Half years
i= (6%/2) 3.00% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.553676 * $15,00,000 = $8,30,514
Interest (Annuity) [$1,500,000*6%*6/12] 14.877475 * $45,000 = $6,69,486
Price of Bonds $15,00,000
Bond Amortization Schedule
Date Cash interest Interest Expense Change in Carrying Value Carrying value
01-Jan-21 $15,00,000
30-Jun-21 $45,000 $45,000 $0 $15,00,000
31-Dec-21 $45,000 $45,000 $0 $15,00,000

Solution 2:

Chart Values are based on:
n= (10 Years*2) 20 Half years
i= (7%/2) 3.50% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.502566 * $15,00,000 = $7,53,849
Interest (Annuity) [$1,500,000*6%*6/12] 14.212403 * $45,000 = $6,39,558
Price of Bonds $13,93,407
Bond Amortization Schedule
Date Cash interest Interest Expense Change in Carrying Value Carrying value
01-Jan-21 $13,93,407
30-Jun-21 $45,000 $48,769 $3,769 $13,97,176
31-Dec-21 $45,000 $48,901 $3,901 $14,01,077

Solution 3:

Chart Values are based on:
n= (10 Years*2) 20 Half years
i= (5%/2) 2.50% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.610271 * $15,00,000 = $9,15,406
Interest (Annuity) [$1,500,000*6%*6/12] 15.589162 * $45,000 = $7,01,512
Price of Bonds $16,16,919
Bond Amortization Schedule
Date Cash interest Interest Expense Change in Carrying Value Carrying value
01-Jan-21 $16,16,919
30-Jun-21 $45,000 $40,423 -$4,577 $16,12,342
31-Dec-21 $45,000 $40,309 -$4,691 $16,07,650